Find the mean of the probability distribution of a random variable X which can take only the values 2, 4, 5, and 9 given the respective probabilities; 9/20, 1/20, 1/5, 3/10. 


1
Expert's answer
2022-03-07T17:13:03-0500

"E(X)={\frac 9 {20}}*2+{\frac 1 {20}}*4+{\frac 1 5}*5+{\frac 3 {10}}*9={\frac {96} {20}}=4.8"